OpenStudy (anonymous):

My physic teacher asks weird questions...can someone please help me convert eyeblinks to armlength? 50 m/s to arm length, eye blinks 59= arm length 61= eye blinks?

OpenStudy (aum):

Weird question indeed. But don't think of the usual meaning of eye blinks and arm length but instead treat them as units in some weird place. 59 eye blinks = 61 arm lengths x eye blinks = y arm lengths Proportions problem. x / 59 = y / 61 y = 61/59 * x Put whatever value for x and calculate y. x eye blinks = y arm lengths.